OpenStudy (anonymous):

If f(x) = 2x − 4 and g(x) = x2 − 4x, find g(2)

geerky42 (geerky42):

Just replace x to 2, like this: \[\large g(x) = x^2-4x\\\large g(2) = (2)^2-4(2)\]

OpenStudy (anonymous):

which would be?

geerky42 (geerky42):

Good question. I'm sure you can calculate \((2)^2-4(2)\)...

geerky42 (geerky42):

\(\Large 2^2-4\cdot2 ~\rightarrow~ 2\cdot2-4\cdot2\)

OpenStudy (anonymous):

0

OpenStudy (anonymous):

or -4. right?